OpenStudy (anonymous):

what is period, amplitude, and range of y=7+3sin2(x-2.5)

zepdrix (zepdrix):

Hey Michelle, \(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

So here is the general form, \[\Large\sf y=\color{orangered}{a}\sin \color{#3366CF}{b}(x-\color{#CC0033}{c})+\color{#A57F02 }{d}\] \(\Large\sf\color{orangered}{a}\) gives us the amplitude, \(\Large\sf 2\pi/\color{#3366CF}{b}\) gives us the period, \(\Large\sf \color{#CC0033}{c}\) is a horizontal shift to the right, and \(\Large\sf \color{#A57F02}{d}\) is a vertical shift upward. Applying this to our problem,\[\Large\sf y=\color{orangered}{3}\sin \color{#3366CF}{2}(x-\color{#CC0033}{2.5})+\color{#A57F02 }{7}\]

zepdrix (zepdrix):

Oh we also need to deal with the range right? Hmm let's see.

zepdrix (zepdrix):

Normally sine goes up to 1, and down to -1. That's when it starts at zero and has an amplitude of 1. So for our problem we instead start at 7, and our amplitude tells us we can go up 3 from there and down 3 from there. So our function has a range from 4 to 10. Does that make sense? :o

OpenStudy (anonymous):

yess it does thank you. i never understood that you needed to put the 7 after everything else :) thanks a lot!!